OpenStudy (anonymous):

write the augmented matrix for the system linear equation 4x-5y=33 -x+5y=-27

OpenStudy (anonymous):

\[ \left[ \begin{array}{ccc} 4 & -5 & 33 \\ -1 & 5 & -27 \end{array} \right]\] There should be a vertical line in between the second and third columns, I forget how to typeset that.

OpenStudy (anonymous):

thnx

OpenStudy (amistre64):

just put in a &|&

OpenStudy (anonymous):

Oh, that makes sense. :P Thanks

OpenStudy (amistre64):

yw \left[ \begin{array}{ccc} 4 & -5 & 33 \\ -1 & 5 & -27 \end{array} \right] this can be simplified ... begin{vmatrix}4 & -5 &|& 33 \\ -1 & 5 &|& -27 end{vmatrix} to get \[\begin{vmatrix} 4 & -5 &|& 33 \\ -1 & 5 &|& -27 \end{vmatrix}\]

OpenStudy (anonymous):

Oh cool. I'm really new to TeX. Trying to study up on it.

OpenStudy (amistre64):

soo hard to try to relate latex in a latex parser lol good luck with it, its can be a pain

OpenStudy (anonymous):

Yeah, it's definitely a pain, but the results are pretty cool :D